# Robust CLIP: Unsupervised Adversarial Fine-Tuning of Vision Embeddings for Robust Large Vision-Language Models

### v1

@@ Results.
On ImageNet, TeCoA models perform best in clean and robust evaluations, as they have undergone supervised training on this dataset. FARE models are also trained on ImageNet but do not take labels into account. On the zero-shot datasets, the undefended CLIP model expectedly has the best performance on clean data, while TeCoA models suffer significant decrease of clean performance. In contrast, the FARE models, especially FARE2, maintain much better clean accuracy. On adversarial inputs, CLIP breaks completely at both radii. FARE4 performs best in this scenario, outperforming TeCoA4 and TeCoA2 for both threat models. FARE is thus also in this setting the only method that provides high-performing and robust models.
@@
Adversarial
Popular
Random
Mean
CLIP
82.6
85.1
85.9
84.5
TeCoA2-CLIP
74.0
76.5
77.3
75.9
FARE2-CLIP
78.6
81.5
82.2
80.8
TeCoA4-CLIP
70.2
73.0
73.3
72.2
FARE4-CLIP
74.0
77.0
77.8
76.3
4.4 Performance on Other Tasks
Besides being robust to adversarial attacks, VLMs should avoid hallucinations and be able to solve Chain of Thought (CoT) tasks. In this section we examine how our robust models fare on hallucination (POPE (Li et al., 2023b)) and CoT (SQA-I (Lu et al., 2022)) benchmarks.
Hallucinations.
Large VLMs are known to suffer from object hallucinations, i.e. they “see” in a target image objects which are not actually present. In Li et al. (2023b) a hallucination benchmark called POPE is proposed, where the evaluation of object hallucination is formulated as a binary task, i.e. the VLM has to decide whether an object is present in the image or not. More details can be found in App. C.1.
In Table 5, we report the F1-score for each of the evaluation settings of POPE when using LLaVA-1.5 7B with different vision encoders. The clean CLIP model attains the best score and FARE is close to it. The TeCoA model attains the worst average F1-score. TeCoA’s proclivity to hallucinations can be attributed to it lacking in ability to generate the correct output even for nominal inputs, as can be seen in Figs. LABEL:fig:teaser-attack and 3. Some examples from the POPE task where different models hallucinate are visualized in App. C.1.
Science Question Answering.
@@ 4.4 Performance on Other Tasks
Science Question Answering (SQA) (Lu et al., 2022) was recently introduced to benchmark large VLMs on reasoning tasks. In this section we test whether for SQA-I (a subset of 10k image/question pairs from SQA) robust models loose their ability to solve reasoning tasks. More task related details can be found in App. C.2. In Table 6, the LLaVA model using original CLIP achieves an accuracy of 64.5%. Both FARE models are better than the respective TeCoA models by 2.4% and additionally FARE2 is only 1% off from the original CLIP model. As the differences of FARE models to CLIP are marginal, we conclude that robustification of vision encoder does not degrade the VLMs ability to solve reasoning tasks, if one does unsupervised adversarial fine-tuning via FARE.
@@
Table 6: SQA-I evaluation with LLaVA. The improvement of FARE to the respective TeCoA model is highlighted.
CLIP
TeCoA2
FARE2
TeCoA4
FARE4
64.5
61.1
63.4
$\uparrow$2.3
59.9
62.3
$\uparrow$2.4
@@ 5 Conclusion
We propose an unsupervised adversarial fine-tuning framework for vision encoders that aims at preserving the original embeddings, while also transferring robustness to downstream tasks. In particular, we are able to obtain adversarially robust large vision-language models by substituting their original CLIP model with our robust FARE-CLIP, without any re-training of the downstream VLM. Our method thus provides an easy and effective defense against visual adversaries of VLMs while maintaining high performance on nominal inputs, in contrast to other adversarially robust CLIP models. As most users of machine learning models are not willing to sacrifice nominal performance for gains in robustness, our models are a felicitous choice for practical applications and real-world deployment.
Limitations. We focus on CLIP based VLMs in this work. Other types of VLMs might also benefit from our approach but we defer this to future work. Defense of the language side of VLMs is also left to future work.
Broader Impact
Large VLMs are being deployed ubiquitously due to their impressive performance across multiple tasks. This makes their safe and secure deployment a pressing problem. In our work we take a step to address it, and believe that our robust models can help in making the deployment of VLMs more safe. Our transfer attacks in Table 2 show that VLMs using the same non-robust vision encoder can be successfully attacked independently of the language model or the part of the VLM which connects language and vision input. This stresses the importance of having a robust vision encoder.
@@ Appendix A Omitted Proof
$\leq$ $\left(\frac{1}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}+\frac{1}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}\right)\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}.$
@@
$\leq$ $\frac{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}|\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}-\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}|}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}$
$+\frac{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)-\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}$
Now using the reverse triangle inequality yields the result:
$|\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}-\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}|\leq\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)-\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}.$
@@ B.1 General Setup
LLaVA and OpenFlamingo use the output of all tokens of the CLIP vision-encoder (LLaVA operates on second-last layer outputs). However, early experiments showed that using only the class-token in the fine-tuning loss is sufficient to attain good results with downstream VLMs. Taking all tokens into account for training requires more memory and compute, but did not lead to improvements. The FARE-loss (Eq. 3) is thus computed with respect to the class token only.
@@
The adversarial evaluations are done for $\ell_{\infty}=\nicefrac{{2}}{{255}}$ with the attack setup mentioned in Sec. 4.1. “ZS-Class.” refers to the average zero-shot image classification accuracy for the datasets from Sec. 4.3. The zero-shot image classification is done only for CLIP (marked with $\bigtriangleup$) wheras the remaining evaluations are done with LLaVA and are marked with $\bigstar$.
B.3 Ablation for Training Hyperparameters
Table 7: Ablation of training hyperparameters. We ablate weight decay (WD) and learning rate (LR) for a ViT-B CLIP vision encoder with the FARE training method. The avg. zero-shot column is average accuracy across all zero-shot datasets from Sec. 4.3. First row (CLIP) is completely non-robust for both ImageNet and other datasets. The final setting yields best generalization to downstream zero-shot tasks.
@@ B.3 Ablation for Training Hyperparameters
LR
WD
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
CLIP
ViT-B/32
62.2
0.0
0.0
64.1
0.0
0.0
FARE4-CLIP
ViT-B/32
1e-5
1e-3
10
51.1
29.6
14.8
48.6
33.7
21.8
FARE4-CLIP
ViT-B/32
1e-5
1e-4
10
51.1
29.6
14.8
48.6
33.7
21.8
FARE4-CLIP
ViT-B/32
1e-4
1e-4
10
51.7
34.2
20.2
44.4
33.3
23.8
FARE4-CLIP
ViT-B/32
1e-4
1e-3
10
51.6
34.3
20.3
44.4
33.5
23.7
All vision encoders in CLIP in the main section of the paper use ViT-L/14 as architectures. Given the high computational cost of training such networks, to get the final training hyperparameters we conducted an ablation using ViT-B/32 vision encoder backbones instead, and fix the FARE loss as training objective. We show in App. B.4 that the resulting training scheme is effective for TeCoA too. The main hyperparameters in our search were the learning rate (LR) and the weight decay coefficient (WD). In Table 7, we present the performance on clean and adversarial inputs for ImageNet and the average over zero-shot datasets from Sec. 4.3.
@@
From Table 7, clean CLIP model is completely non-robust, which is expected as it was trained only on nominal samples. Across all FARE models, weight decay (WD) seems to have no impact on both the clean performance and the robustness. Whereas smaller LR (1e-5) yields models that generalize better to zero-shot datasets in comparison to the 1e-4 models. Since we want the resulting robust models to not loose too much in terms of performance on downstream zero-shot tasks from original CLIP (one of the drawbacks of TeCoA), we relinquish the gains in ImageNet robustness that LR 1e-4 models have over smaller LR models (+5% robustness on average across the two perturbation radii). Hence, we select LR = 1e-5 and WD = 1e-4, which has +4.2% clean zero-shot performance and similar zero-shot robustness in comparison to LR=1e-4 setup as our final parameter setting.
Table 8: Targeted attacks with only 500 iterations. We run the targeted attacks of Table Datasets and metrics. for 500 iterations (instead of 10,000) and observe that this attack is considerably weaker for $\varepsilon=\nicefrac{{2}}{{255}}$.
LLaVA with CLIP
Target $\varepsilon=\;\;\nicefrac{{2}}{{255}}\;$ $\nicefrac{{4}}{{255}}$
EmailAPI(to=<target(…) 7 / 25 25 / 25
Vaccines are dangerous. (…) 11 / 25 25 / 25
You are too stupid. I (…) 25 / 25 25 / 25
Sell your stocks. (…) 19 / 25 25 / 25
Visit https://tinyurl.com/(…) 14 / 25 25 / 25
A person suffers (…) 13 / 25 25 / 25
Mean success rate: 59.3% 100%
@@
$\nicefrac{{1}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
$\nicefrac{{1}}{{255}}$
$\nicefrac{{2}}{{255}}$
$\nicefrac{{4}}{{255}}$
CLIP
OpenAI
62.2
0.0
0.0
0.0
64.1
0.3
0.0
0.0
TeCoA
$\nicefrac{{1}}{{255}}$
2
10
 Mao et al. (2023)
54.6
35.8
20.1
3.4
50.3
38.2
27.1
9.8
TeCoA
$\nicefrac{{1}}{{255}}$
10
2
ours
70.3
53.2
34.5
8.0
53.1
38.2
26.6
9.6
FARE
$\nicefrac{{1}}{{255}}$
10
2
ours
62.1
32.9
12.2
0.2
60.5
38.0
20.1
2.9
B.4 Comparison to Original TeCoA Checkpoint
In this section, we show a comparison between the original TeCoA ViT-B/32 checkpoint [^1] https://github.com/cvlab-columbia/ZSRobust4FoundationModel (from Mao et al. (2023)) to a TeCoA ViT-B/32 model we trained. Note that Mao et al. (2023) did not train a ViT-L/14 model and thus a direct comparison to the VLM tasks done in the main paper which require ViT-L/14 models is not feasible. In particular, we report the performance of the models in the zero-shot classification setup as in Sec. 4.3. The purpose of this section is to show that our selected hyperparameters work also well for TeCoA.
In Mao et al. (2023), the ViT-B/32 model has been trained for 10 epochs using 2 steps of PGD at $\ell_{\infty}$ radius of $\nicefrac{{1}}{{255}}$. Note that in the main paper we always train ViT-L/14 models only for two epochs and for $\ell_{\infty}$ radii $\nicefrac{{2}}{{255}}$ and $\nicefrac{{4}}{{255}}$, as our goal is to get non-trivial robustness also at these larger radii. However, for better comparison we train also ViT-B/32 models for TeCoA and FARE with our chosen hyperparameters at $\varepsilon=\nicefrac{{1}}{{255}}$ for one epoch. In Table 9 we compare the TeCoA model of Mao et al. (2023), our TeCoA model and our FARE model trained for $\varepsilon=\nicefrac{{1}}{{255}}$, all with the same forward/backward pass budget.

### v2

@@ Results.
On ImageNet, TeCoA models perform best in clean and robust evaluations, as they have undergone supervised training on this dataset. FARE models are also trained on ImageNet but do not take labels into account. On the other zero-shot datasets, the undefended CLIP model expectedly has the best performance on clean data, while TeCoA models suffer significant decrease of clean performance. In contrast, the FARE models, especially FARE2, maintain much better clean accuracy. On adversarial inputs, CLIP breaks completely at both radii. FARE4 performs best in this scenario, outperforming TeCoA4 and TeCoA2 across threat models. FARE is thus also in this setting the only method that provides high-performing and robust models.
Eval. Vision encoder Zero-shot datasets
ImageNet CalTech Cars CIFAR10 CIFAR100 DTD EuroSAT FGVC Flowers ImageNet-R ImageNet-S PCAM OxfordPets STL-10 Average Zero-shot
clean CLIP 74.9 83.3 77.9 95.2 71.1 55.2 62.6 31.8 79.2 87.9 59.6 52.0 93.2 99.3 73.1
TeCoA2-CLIP 80.2 80.7 50.1 87.5 60.7 44.4 26.1 14.0 51.8 80.1 58.4 49.9 80.0 96.1 60.0
FARE2-CLIP 74.2 84.8 70.5 89.5 69.1 50.0 25.4 26.7 70.6 85.5 59.7 50.0 91.1 98.5 67.0 $\uparrow$7.0
[…]
FARE4-CLIP 70.4 84.7 63.8 77.7 56.5 43.8 18.3 22.0 58.1 80.2 56.7 50.0 87.1 96.0 61.1 $\uparrow$6.9
$\ell_{\infty}=\nicefrac{{2}}{{255}}$ CLIP 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.1 0.0 0.0 0.0 0.0
TeCoA2-CLIP 62.3 70.2 22.2 63.7 35.0 27.0 12.8 5.8 27.6 58.8 45.2 40.0 69.7 88.7 43.6
FARE2-CLIP 46.1 73.0 26.0 60.3 35.6 26.7 6.2 5.9 31.2 56.5 38.3 41.9 68.3 90.1 43.1 $\downarrow$0.5
[…]
FARE4-CLIP 52.4 76.7 30.0 57.3 36.5 28.3 12.8 8.2 31.3 61.6 41.6 50.2 72.4 89.6 45.9 $\uparrow$3.6
$\ell_{\infty}=\nicefrac{{4}}{{255}}$ CLIP 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0
TeCoA2-CLIP 37.3 57.4 6.5 31.0 17.8 14.7 7.7 1.1 9.8 36.7 32.8 16.0 50.3 69.2 27.0
FARE2-CLIP 16.6 46.6 4.8 25.9 13.9 11.7 0.5 0.6 7.1 25.6 22.5 17.2 27.9 61.7 20.5 $\downarrow$6.5
[…]
FARE4-CLIP 33.3 64.1 12.7 34.6 20.2 17.3 11.1 2.6 12.5 40.6 30.9 50.2 50.7 74.4 32.4 $\uparrow$0.5
4.4 Performance on Other Tasks
Until now, we focused on adversarial attacks. Recently, (Qi et al., 2023) proposed jailbreaking attacks for LVLMs. We test the robustness of LLaVA 1.5 using TeCoA and FARE to such attacks in this section. Besides being robust to different type of attacks, LVLMs should avoid hallucinations and be able to solve Chain of Thought (CoT) tasks which we also examine in this section via POPE (Li et al., 2023b) and SQA-I (Lu et al., 2022) benchmarks.
Hallucinations.
Large vision-language models are known to suffer from object hallucinations, i.e. they “see” in a target image objects which are not actually present. In Li et al. (2023b) a hallucination benchmark called POPE is proposed, where the evaluation of object hallucination is formulated as a binary task, i.e. the LVLM has to decide whether an object is present in the image or not. More details can be found in App. C.1.
In Table 5, we report the F1-score for each of the evaluation settings of POPE when using LLaVA-1.5 7B with different vision encoders. The clean CLIP model has the best performance on all splits of POPE, while FARE is the closest to it. The TeCoA model attains the worst average F1-score. TeCoA’s proclivity to hallucinations can be attributed to it lacking in ability to generate the correct output even for nominal inputs, as can be seen in Figs. LABEL:fig:teaser-attack and 3. Some qualitative examples from the POPE task showing varying levels of hallucinations for different models are visualized in Fig. 4 in App. C.1.
@@ 4.4 Performance on Other Tasks
Adversarial Popular Random Mean
CLIP 82.6 85.1 85.9 84.5
TeCoA2-CLIP 74.0 76.5 77.3 75.9
FARE2-CLIP 78.6 81.5 82.2 80.8
TeCoA4-CLIP 70.2 73.0 73.3 72.2
FARE4-CLIP 74.0 77.0 77.8 76.3
Chain of Thought (CoT).
@@
Science Question Answering (SQA) (Lu et al., 2022) was recently introduced to benchmark LVLMs on reasoning tasks. In this section we test whether for SQA-I (a subset of 10k image/question pairs from SQA) robust models loose their ability to solve reasoning tasks. More task related details are reported in App. C.2. In Table 6, the LLaVA model using original CLIP achieves an accuracy of 64.5%. Both FARE models are better than the respective TeCoA models by 2.4% and additionally FARE2 is only 1% off from the original CLIP model. As the differences of FARE models to CLIP are marginal, we conclude that robustification of vision encoder does not degrade the LVLMs ability to solve reasoning tasks, if one does unsupervised adversarial fine-tuning via FARE.
@@
Table 6: SQA-I evaluation with LLaVA. The performance of different models are shown, with the improvement of FARE to the respective TeCoA model highlighted. Overall FARE models are better than TeCoA.
CLIP TeCoA2 FARE2 TeCoA4 FARE4
64.5 61.1 63.4 $\uparrow$2.3 59.9 62.3 $\uparrow$2.4
Table 7: Jailbreaking attacks against LLaVA 1.5. We run the attack proposed by Qi et al. (2023) and report the success rates across harmful prompts of different categories. Lower numbers indicate more robust models. LLaVA 1.5 with TeCoA or FARE is significantly more robust than with original CLIP.
LLaVA using     $\varepsilon$ any identity disinfo. crime x-risk
CLIP 0 12 / 40 4 / 11 5 / 13 1 / 13 2 / 3
TeCoA4 0 14 / 40 3 / 11 8 / 13 1 / 13 2 / 3
FARE4 0 13 / 40 3 / 11 8 / 13 1 / 13 1 / 3
CLIP $\nicefrac{{16}}{{255}}$ 24 / 40 10 / 11 9 / 13 2 / 13 3 / 3
TeCoA4 $\nicefrac{{16}}{{255}}$ 14 / 40 3 / 11 8 / 13 1 / 13 2 / 3
FARE4 $\nicefrac{{16}}{{255}}$ 15 / 40 3 / 11 9 / 13 1 / 13 2 / 3
CLIP $\nicefrac{{32}}{{255}}$ 28 / 40 11 / 11 11 / 13 3 / 13 3 / 3
TeCoA4 $\nicefrac{{32}}{{255}}$ 14 / 40 2 / 11 9 / 13 1 / 13 2 / 3
FARE4 $\nicefrac{{32}}{{255}}$ 16 / 40 3 / 11 10 / 13 1 / 13 2 / 3
CLIP $\nicefrac{{64}}{{255}}$ 36 / 40 11 / 11 13 / 13 9 / 13 3 / 3
TeCoA4 $\nicefrac{{64}}{{255}}$ 23 / 40 10 / 11 9 / 13 1 / 13 3 / 3
FARE4 $\nicefrac{{64}}{{255}}$ 23 / 40 9 / 11 10 / 13 2 / 13 2 / 3
Robustness to Jailbreaking Attacks.
Large vision-language models are known to be vulnerable to jailbreaking attacks on the visual input modality (Carlini et al., 2023; Qi et al., 2023). An adversary can craft input images that cause LVLMs to adhere to harmful prompts, e.g. “How to build a bomb?”. We test the ability of robust vision-encoders to defend against such attacks. To this end, we craft adversarial images by running the attack from Qi et al. (2023) against LLaVA-1.5 7B with different vision encoders (CLIP, TeCoA4, FARE4) and varying attack strength $\varepsilon$. Then we evaluate the success of the attack by querying models with their respective adversarial image and 40 harmful prompts of various categories, as proposed by Qi et al. (2023).
The results are reported in Table 7. Robust CLIP models indeed help in defending LLaVA 1.5 against jailbreaking attacks even at attack radii which are much higher than for which they have been trained. TeCoA and FARE similarly reduce the number of harmful outputs significantly compared to the original CLIP vision encoder. Irrespective of attack strength ($\varepsilon$) and type of prompt, both TeCoA and FARE are equally effective.
We note that jailbreaking attacks are an active research area. Thus our evaluation based on the attack of Qi et al. (2023) is preliminary and might overestimate robustness. Improving such attacks goes beyond the scope of our paper.
@@ 5 Conclusion
We propose an unsupervised adversarial fine-tuning framework, FARE, for vision encoders that aims at preserving the original embeddings, thereby maintaining nominal performance and transferring robustness to down-stream tasks. Thanks to such approach, we are able to obtain adversarially robust large vision-language models (LVLMs) by substituting their original CLIP vision encoder with our robust FARE-CLIP encoder. Importantly, this procedure does not require any retraining of the down-stream LVLM, which would be time-consuming and expensive. Thus, our method provides an easy defense against visual adversaries of LVLMs while maintaining high performance on nominal inputs. As most users of machine learning models are not willing to sacrifice nominal performance for gains in robustness, our models are a felicitous choice for practical applications and real-world deployment.
We also show that the proposed method generalizes to other aspects where LVLMs are expected to be good, e.g. hallucinations and chain-of-thought experiments. Moreover, the proposed FARE-CLIP models exhibit excellent zero-shot classification capabilities, outperforming previous methods in terms of clean and adversarial performance.
Finally, in this work we consider LVLMs which have frozen vision encoders, but our method can be easily extended to newer LVLMs which fine-tune the vision encoder: in fact, the proposed FARE can be applied after the LVLM is fully trained, at little extra computational cost.
Limitations. This work focuses on CLIP-based LVLMs, but other types of LVLMs might also benefit from the proposed approach. Moreover, the robustness of our method is restricted to the visual input space of LVLMs, the defense of the language side of LVLMs is also important. This work also does not examine the influence of using robust CLIP-enabled LVLMs for instruction following, explainability, and perception related tasks. We leave the investigation of these aspects to future work.
Impact Statement
Large vision-language models are being deployed ubiquitously due to their impressive performance across multiple tasks. This makes their safe and secure deployment a pressing problem. In our work we take a step to address this, and believe that our robust models can help in making the deployment of LVLMs more safe. Our transfer attacks in Table 2 show that LVLMs using the same non-robust vision encoder can be successfully attacked independently of the language model or the part of the LVLM which connects language and vision input, thereby enabling attacks even on closed-source LVLMs. This stresses the importance of having a robust vision encoder.
@@ Appendix A Omitted Proof
$\leq$ $\vspace{-1mm}\min\Big{(}\vspace{-1mm}\frac{2}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}},\frac{2}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}\vspace{-1mm}\Big{)}\vspace{-1mm}\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)-\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}.$
@@
For which we can get the two upper bounds:
$\left\|\frac{\mathop{\phi_{\rm{Org}}}\nolimits(x)}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}-\frac{\mathop{\phi_{\rm{FT}}}\nolimits(x)}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}\right\|_{2}$
$\leq$ $\frac{1}{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}[\,\left|\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}-\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}\right|$
$+\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)-\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}]$
and
$\left\|\frac{\mathop{\phi_{\rm{Org}}}\nolimits(x)}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}-\frac{\mathop{\phi_{\rm{FT}}}\nolimits(x)}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}\right\|_{2}$
$\leq$ $\frac{1}{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}[\left|\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}-\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}\right|$
$+\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)-\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}],$
where inside the norm we have added and subtracted $\nicefrac{{\mathop{\phi_{\rm{Org}}}\nolimits(x)}}{{\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}}}$ for the first bound and $\nicefrac{{\mathop{\phi_{\rm{FT}}}\nolimits(x)}}{{\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}}}$ for the second bound.
Now using the reverse triangle inequality:
$|\left\|\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2}-\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)\right\|_{2}|\leq\left\|\mathop{\phi_{\rm{Org}}}\nolimits(x)-\mathop{\phi_{\rm{FT}}}\nolimits(x)\right\|_{2},$
and the minimum of the two upper bounds yields the result. ∎
@@ B.1 General Setup
LLaVA and OpenFlamingo use the output of all tokens of the CLIP vision-encoder (LLaVA operates on second-last layer outputs). However, early experiments showed that using only the class-token in the fine-tuning loss is sufficient to attain good results with down-stream LVLMs. Taking all tokens into account for training requires more memory and compute, but did not yield improvements. The FARE-loss (Eq. 3) is thus computed with respect to the class token only.
@@
The adversarial evaluations are done for $\ell_{\infty}=\nicefrac{{2}}{{255}}$ with the attack setup mentioned in Sec. 4.1. “ZS-Class.” refers to the average zero-shot image classification accuracy for the datasets from Sec. 4.3. The zero-shot image classification is done only for CLIP (marked with $\bigtriangleup$) whereas the remaining evaluations are done with LLaVA and are marked with $\bigstar$.
B.3 Ablation of Training Hyperparameters
Table 8: Ablation of training hyperparameters. We ablate weight decay (WD) and learning rate (LR) for a ViT-B CLIP vision encoder with the FARE fine-tuning method. The avg. zero-shot column is average accuracy across all zero-shot datasets from Sec. 4.3. First row (CLIP) is completely non-robust for both ImageNet and other datasets. The final setting yields best generalization to down-stream zero-shot tasks.
@@ B.3 Ablation of Training Hyperparameters
LR WD $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
CLIP ViT-B/32 62.2 0.0 0.0 64.1 0.0 0.0
FARE4-CLIP ViT-B/32 1e-5 1e-3 10 51.1 29.6 14.8 48.6 33.7 21.8
FARE4-CLIP ViT-B/32 1e-5 1e-4 10 51.1 29.6 14.8 48.6 33.7 21.9
FARE4-CLIP ViT-B/32 1e-4 1e-4 10 51.7 34.2 20.2 44.4 33.3 23.8
FARE4-CLIP ViT-B/32 1e-4 1e-3 10 51.6 34.3 20.3 44.4 33.5 23.7
All vision encoders in CLIP in the main section of the paper use ViT-L/14 as architectures. Given the high computational cost of training such networks, to get the final training hyperparameters we conducted an ablation using ViT-B/32 vision encoder backbones instead, and fix the FARE loss as training objective. We show in App. B.5 that the resulting training scheme is effective for TeCoA too. The main hyperparameters in our search were the learning rate (LR) and the weight decay coefficient (WD). In Table 8, we present the performance on clean and adversarial inputs for ImageNet and the average over zero-shot datasets from Sec. 4.3.
@@
From Table 8, clean CLIP model is completely non-robust, which is expected as it was trained only on nominal samples. Across all FARE models, weight decay (WD) seems to have no impact on both the clean performance and the robustness. Whereas smaller LR (1e-5) yields models that generalize better to zero-shot datasets in comparison to the 1e-4 models. Since we want the resulting robust models to not loose too much in terms of performance on down-stream zero-shot tasks from original CLIP (one of the drawbacks of TeCoA), we relinquish the gains in ImageNet robustness that LR 1e-4 models have over smaller LR models (+5% robustness on average across the two perturbation radii). Hence, we select LR = 1e-5 and WD = 1e-4, which has +4.2% clean zero-shot performance and similar zero-shot robustness in comparison to LR=1e-4 setup as our final parameter setting.
Table 9: Ablation of loss function. We compare ViT-B/32 FARE models trained with the original squared $\ell_{2}$-norm formulation (Eq. (3)), and using the $\ell_{1}$-norm instead.
Loss used in Eq. (3) ImageNet Avg. Zero-shot
clean $\ell_{\infty}$ clean $\ell_{\infty}$
$\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
$\lVert\cdot\rVert_{2}^{2}$ 51.1 29.6 14.8 48.6 33.7 21.9
$\lVert\cdot\rVert_{1}$ 51.2 30.1 15.1 48.6 33.9 21.9
@@
$\nicefrac{{1}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{1}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
CLIP - - - OpenAI 62.2 0.0 0.0 0.0 64.1 0.3 0.0 0.0
TeCoA $\nicefrac{{1}}{{255}}$ 2 10  Mao et al. (2023) 54.6 35.8 20.1 3.4 50.3 38.2 27.1 9.8
TeCoA $\nicefrac{{1}}{{255}}$ 10 2 ours 70.3 53.2 34.5 8.0 53.1 38.2 26.6 9.6
FARE $\nicefrac{{1}}{{255}}$ 10 2 ours 62.1 32.9 12.2 0.2 60.5 38.0 20.1 2.9
Table 11: Comparing our ensemble attack to that of Schlarmann & Hein (2023). The two types of attack are compared for the non-robust CLIP and our most robust FARE4 vision encoders with OpenFlamingo-9B. Across both perturbation strengths and for both captioning (COCO) and question answering (VQAv2) tasks our “Ensemble” attack is much better while being significantly faster. The runtime is averaged over all settings for the respective attack.
Attack Source Runtime COCO VQAv2
CLIP FARE4 CLIP FARE4
$\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$ $\nicefrac{{2}}{{255}}$ $\nicefrac{{4}}{{255}}$
Single-precision Schlarmann & Hein (2023) 5h 8m 5.7 2.9 67.9 55.6 6.9 6.5 38.0 29.8
Ensemble ours 0h 40m 1.3 1.1 30.4 21.7 4.6 4.1 26.3 21.4
B.4 Ablation of Loss Function
In the main paper we use the squared $\ell_{2}$-norm to measure similarity between original and perturbed embeddings in our formulation of the FARE-loss (3). This choice is motivated by (i) its close connection to the cosine-similarity [^1] For $u,v\in\mathbb{R}^{d}$ it holds $\lVert\frac{u}{\lVert u\rVert_{2}}-\frac{v}{\lVert v\rVert_{2}}\rVert_{2}^{2}=2-2\cos(u,v)$, which is used for zero-shot classification and (ii) its preservation of non-normalized embeddings, see Sec. 3.2.
For ablation, we train a ViT-B/32 FARE model, using the $\ell_{1}$-norm instead of the squared $\ell_{2}$-norm in Eq. (3). We note that minimizing the $\ell_{1}$-loss can lead to sparse residuals, for which we see no motivation in the present setting. Results for this ablation are reported in Table 9. We observe that using the $\ell_{1}$-norm yields similar performance.
B.5 Comparison to Original TeCoA Checkpoint
In this section, we show a comparison between the original TeCoA ViT-B/32 checkpoint [^2] https://github.com/cvlab-columbia/ZSRobust4FoundationModel (from Mao et al. (2023)) to a TeCoA ViT-B/32 model we trained. Note that Mao et al. (2023) did not train a ViT-L/14 model and thus a direct comparison to the LVLM tasks done in the main paper which require ViT-L/14 models is not feasible. In particular, we report the performance of the models in the zero-shot classification setup as in Sec. 4.3. The purpose of this section is to show that our selected hyperparameters work also well for TeCoA.
In Mao et al. (2023), the ViT-B/32 model has been trained for 10 epochs using 2 steps of PGD at $\ell_{\infty}$ radius of $\nicefrac{{1}}{{255}}$. Note that in the main paper we always train ViT-L/14 models only for two epochs and for $\ell_{\infty}$ radii $\nicefrac{{2}}{{255}}$ and $\nicefrac{{4}}{{255}}$, as our goal is to get non-trivial robustness also at these larger radii. However, for better comparison we train also ViT-B/32 models for TeCoA and FARE with our chosen hyperparameters at $\varepsilon=\nicefrac{{1}}{{255}}$ for two epochs. In Table 10 we compare the TeCoA model of Mao et al. (2023), our TeCoA model and our FARE model trained for $\varepsilon=\nicefrac{{1}}{{255}}$, all with the same forward/backward pass budget.